makaylab:

Tracy deposited $59 into a bank account that earned 3.5% simple interest each year. No money was deposited into or withdrawn from the account. How much money was in the account after ​ 112 ​ years? Round your answer to the nearest cent.

dude:

Welcome to QuestionCove! The simple interest formula could help you solve this equation \(A= P(1+rt)\) A = total amount P = principal(starting) amount r = rate t = time Do you know how to use this equation?

makaylab:

idk

dude:

Okay lets start with the basics What numbers are they giving us?

makaylab:

59 3.5% and 112

makaylab:

now what do i do

dude:

Yeah Do you know what each stands for?

makaylab:

no

dude:

Hmm What is the unit on 112?

makaylab:

idk *cries inside*

dude:

Just review the question :) `How much money was in the account after ​112 ​ years?`

makaylab:

idk

makaylab:

that is what i am trying to figure out

dude:

Its nothing out of the ordinary 112 just refers to years

dude:

Yes?

makaylab:

what do you mean tho

dude:

The question just says they want to know about the amount after 112 years 112 years is the time we want to know, right?

makaylab:

yea

dude:

Cool cool 3.5 is given as a percent, also stating it is interest Do you know which of these it would be? \(\color{#0cbb34}{\text{Originally Posted by}}\) A = total amount P = principal(starting) amount r = rate \(\color{#0cbb34}{\text{End of Quote}}\)

makaylab:

ok so what do i do to solve

dude:

Lets hold on I want to make sure you know what each means so that you have an idea on what to do in the future :) I'll ask another thing $59, do you think this is the starting amount, or the rate?

makaylab:

rate

dude:

Careful It would not be rate, it is the starting amount 3.5% is the rate How do find this out? Well usually the question will hint at it by the word `each` followed by a unit of time `each hour`, `each year` are examples of this The question has `earned 3.5% simple interest **each year.**` Low and behold we see each! This is how we know the rate :)

dude:

Now that we know everything we can substitute \(A= P(1+rt)\) We know: t, r, and P Which is: \(A= 59(1+(0.035)(12))\) Can you solve this now?

makaylab:

no well maybe let me try and you can tell me if i am right ok?

dude:

Of course!

b6yhybjfyvgchbujfybj:

can you help me

makaylab:

i can try

b6yhybjfyvgchbujfybj:

\Calculate the correlation coefficient of the data

dude:

@b6yhybjfyvgchbujfybj Hey make a post! I can help you out later, lets keep the replies relevant to the original asker

dude:

@makaylab If you are still not sure yet, you can ask for help

makaylab:

1.42 i think mine is

dude:

Not quite Lets start off slow You'll need a calculator What is \(0.035 \times 12\)?

makaylab:

one sec

makaylab:

0.42

dude:

Good :) What is \(1+0.42\)?

makaylab:

0.43

dude:

Not quite \(1.00\) +\(0.42\)

makaylab:

1.42

dude:

Perfect Now, we lastly do \(59\times 1.42\)

makaylab:

83.78

dude:

Yes! That is your answer Remember units You'll have \($83.78\)

makaylab:

ok thank you so much

dude:

Oh wait I miscounted after D: Forgot it was 112 years, not 12 I apologize for that \(59(1+0.035×112)\) This gives you \($290.28\)

makaylab:

ok thnx